# Uncertainty-Aware Attention Heads: Efficient Unsupervised Uncertainty Quantification for LLMs

### v1

@@ D.1 Experiments Using the ROC-AUC Metric
The results evaluated using the ROC-AUC metric are presented in Table 8. For all generation quality metrics except accuracy, we compute scores by thresholding the original continuous values to obtain discrete versions of the quality metrics. The thresholds were empirically determined as follows: 0.3 for Summ, 0.5 for QA, and 0.85 for MT.
@@ D.2 Comparison with Supervised Methods
We compare our method against several state-of-the-art supervised methods that leverage hidden states or attention weights: Factoscope [20], SAPLMA [1], MIND [41], Sheeps [6], LookBack Lens [8], SATRMD+MSP [45], and TAD [44]. We evaluate these methods in two scenarios: in-domain, where the model is trained directly on the target task, and out-of-domain, where the model is trained on all datasets except one, which is held out for testing. Tables 9 and 10 show the performance of supervised methods in the in-domain and out-of-domain settings respectively.
@@ Appendix E Detailed Experimental Results
The detailed experimental results across each considered dataset are presented in Tables 11, 12, 13 and 14 for Llama-3.1 8b, Qwen-2.5 7b, Gemma-2 9b, and Falcon-3 10b models respectively.
@@ Appendix F Additional Examples
We provide more examples of attention maps, similar to the Figure 1, in Figures 6, 7, 8 and 9. These examples show that the similar patterns exist for several text instances.

### v2

@@ D.1 Experiments Using the ROC-AUC Metric
The results evaluated using the ROC-AUC metric are presented in Table˜8. For all generation quality metrics except accuracy, we compute scores by thresholding the original continuous values to obtain discrete versions of the quality metrics. The thresholds were empirically determined as follows: 0.3 for Summ, 0.5 for QA, and 0.85 for MT.
@@ D.2 Comparison with Supervised Methods
We compare our method against several state-of-the-art supervised methods that leverage hidden states or attention weights: Factoscope [20], SAPLMA [1], MIND [41], Sheeps [6], LookBack Lens [8], SATRMD+MSP [45], and TAD [44]. We evaluate these methods in two scenarios: in-domain, where the model is trained directly on the target task, and out-of-domain, where the model is trained on all datasets except one, which is held out for testing. Tables˜9 and 10 show the performance of supervised methods in the in-domain and out-of-domain settings respectively.
@@ Appendix E Detailed Experimental Results
The detailed experimental results across each considered dataset are presented in Tables˜11, 12, 13 and 14 for Llama-3.1 8b, Qwen-2.5 7b, Gemma-2 9b, and Falcon-3 10b models respectively.
@@ Appendix F Additional Examples
We provide more examples of attention maps, similar to the Figure˜1, in Figures˜6, 7, 8 and 9. These examples show that the similar patterns exist for several text instances.